Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
amagazzu - 15:59 Sunday 18 February 2024 (63309) Print this report
Operator Report - Morning shift

The shift was dedicated to the planned ISC activity, NI Etalon Scan, carried out by the operator. The ITF kept the lock for whole shift. 
At 14:19 UTC, under request of Nardecchia and after the Etalon transient was completed, I changed the Ring Heater setpoints of NE and WE:
NE: 13:2 V
WE 15.3 V
ITF left locked.

Calibrations
7:03 UTC - CALIBRATED_DF_DAILY

Images attached to this report
Search Help
×

Warning

×